OpenStudy (anonymous):

Find: ∫(x5-4x^2+2)dx

OpenStudy (phi):

\[ \int x^5 -4x^2 + 2 \ dx \] integrate term by term. use the "rule" \[ \int x^n \ dx = \frac{1}{n+1} x^{n+1} \]

OpenStudy (anonymous):

Could you solve this one for me? I'm really struggling with homework and want to get it in tonight. We get a completion grade for doing it.

OpenStudy (phi):

which part do you not get? are you saying you can't integrate \[ \int 2 \ dx = 2\int dx\]?

OpenStudy (anonymous):

It's just a blur to me man. Math isn't really my thing and my professor is very hard to understand. Just trying to survive.

OpenStudy (phi):

If you don't know how to integrate \[ \int dx \] then you might try https://www.khanacademy.org/math/calculus/integral-calculus/indefinite_integrals/v/antiderivatives-and-indefinite-integrals